# Pairing by disorder of dopants in a magnetic spin ladder

### v1

@@ Abstract
Figure 3: Steady state. (a) The rms distances $l_{X}$ and $l_{\Delta}$ of the center-of-mass, $X$, and relative, $\Delta$, coordinates of the holes as a function of $J_{\perp}/t$ for indicated values of $J_{\parallel}$ in the steady state. (b) Associated correlation coefficient ${\rm corr}(x_{1},x_{2})=\braket{x_{1}x_{2}}/[\sigma(x_{1})\sigma(x_{2})]$. (c) Relative hole-hole correlator [Eq. (13)] for holes on the same rung ($\Delta=0$) vs $J_{\perp}/t$. (d) Relative hole-hole correlator vs $|\Delta|$ for indicated values of $J_{\perp}/t$ and $J_{\parallel}$.
The dynamical binding – a finite $l_{\Delta}$ – has the exact same origin as the thermal localization phenomenon discovered recently for the motion of single holes Nielsen (2024a, b), by which the separation of the holes is avoided, because the induced interaction $V_{\boldsymbol{\sigma}}(x_{1},x_{2})$ inevitably fluctuates to large values in the relative distance $2|\Delta|=|x_{1}-x_{2}|$, as seen in Fig. 2(b). Therefore, at the very least, they back-reflect once they meet their classical turning points, corresponding to an emergent kind of Anderson localization Anderson (1958) in the presence of strong disorder Lagendijk et al. (2009). The retained localization of the holes – a finite $l_{X}$ – is more subtle, because $V_{{\boldsymbol{\sigma}}}(x,x)$ is restricted to take values between ±$\pm|J_{\parallel}|$ [Fig. 2(c)], such that it vanishes for holes on the same rung in the extreme limit of $J_{\parallel}=0$. Even in this case, however, the holes localize [Fig. 3(b)]. The reason is that there are many ways in which the holes can move to sites $(1,x)$ and $(2,x)$. Along each path, the wave builds up a phase directly depending on the randomly fluctuating induced interaction they meet along the way. The ensuing destructive interference of these pathways corresponds to Anderson localization for weak disorder Lagendijk et al. (2009). Finally, the relative hole-hole correlator in Figs. 3(c) and 3(d) show that the probability of finding holes in close proximity can be several times higher than expected from their marginals. Overall, it is clear that at large $J_{\perp}/t$ the strength of the pairing sensitively depends on $J_{\parallel}$, essentially because the holes localize over very few lattice sites if $J_{\parallel}\sim J_{\perp}$. However, already at intermediate values of $J_{\perp}/t$, the pairing for $J_{\parallel}=J_{\perp}$ is comparable to the case of $J_{\parallel}=0$. I expect this to persist at lower $J_{\perp}/t$.
@@
The predicted results may be feasibly tested in ultracold quantum simulators based on Rydberg-dressed atoms in optical lattices Zeiher et al. (2016). Indeed, such schemes natively realizes spin-specific density-density interactions
${\hat{H}}_{J}=\frac{1}{2}\sum_{{\bf i}\neq{\bf j}}J_{{\bf i}-{\bf j}}\hat{n}_{\uparrow{\bf i}}\hat{n}_{\uparrow{\bf j}}$ (14)
@@
of the internal atomic state $\ket{\uparrow}$, by optically dressing it with a higher-lying Rydberg state. Here, $J_{{\bf i}-{\bf j}}$ takes on a soft core shape Henkel et al. (2010); SM . Moreover, such systems were recently demonstrated to enter the itinerant regime Weckesser et al. (2024), by which the desired $t$-$J$ model may be realized. In particular, the spin model in Eq. (14) is obtained by using a Feshbach resonance to enhance the onsite interaction deep into the Mott-insulating phase. This ensures at most one spin per site and negligible spin superexchange. Precise control of the temperature in such systems is difficult, but the hole dynamics may still be investigated in the following manner Nielsen (2024a). First, initialize holes at $x_{1}=x_{2}=0$ by applying a strong repulsive light field to those sites Ji et al. (2021). Second, rotate all spins from the fully polarized spin sample in the non-interacting $\ket{\downarrow}$ state into the equal superposition state |$\ket{\Psi_{0}}=\prod_{{\bf i}\neq{\bf 0}}\frac{1}{\sqrt{2}}\left[\hat{c}^{\dagger}_{{\bf i}\uparrow}+\hat{c}^{\dagger}_{{\bf i}\downarrow}\right]\ket{0}$. Third, turn off the focused light fields, freeing the holes to move along the ladder Hirthe et al. (2023). Although this at face value seems distinct from the infinite temperature scenario above, the hole-hole correlator takes on the exact same form
$C_{\rm hh}(x_{1},x_{2};\tau)$ $=2^{-N}\sum_{{\boldsymbol{\sigma}}}|a_{\boldsymbol{\sigma}}(x_{1},x_{2};\tau)|^{2},$ (15)
@@
such that the holes move as though the system is at infinite temperature. The polarized Ising form of the interaction is crucial for this exact equivalence, and one should note that spin correlators would presumably differ from their infinite temperature counterpart. The computation now follows the same recipe as before, and the resulting correlation coefficient dynamics [Eq. (12)] and relative correlation function [Eq. (13)] are given in Figs. 4(a) and 4(b), respectively. This is averaged over $N_{S}=1000$ realizations and given in two instances [inset of Fig. 4(a)]: (1) isotropic couplings akin to the $J_{\parallel}=J_{\perp}$ case, and (2) anisotropic couplings where the $30P_{3/2}$ orbital of the Rydberg state is rotated to lie along the rungs using an external magnetic field Zeiher et al. (2016), suppressing coupling in other directions similar to low values of $J_{\parallel}$ SM . Crucially, we observe that the appearing disorder-induced attraction between the holes can be probed for experimentally feasible system sizes and timescales. Indeed, stroboscopic dressing Weckesser et al. (2024) makes $20$ tunneling events feasible, while longer timescales are inhibited by the intrinsic lifetime of the Rydberg state Weckesser (2024). I note that the rung case gives a much more favorable signal, just as was the case for $J_{\parallel}=0$ vs $J_{\parallel}=J_{\perp}$.
@@
Figure 4: Rydberg dressing. (a) Dynamics of the correlation coefficient for isotropic (orange) and strong rung (purple) spin couplings. (b) Associated relative correlator for holes on the same rung, $\Delta=0$. Dashed lines: steady state values. Nearest neighbor rung interaction: $J_{{\bf e}_{y}}=8t$. Insets: orientation of $30P_{3/2}$ Rydberg orbitals enforced by shown magnetic fields ${\bf B}$. System size: $N=21\times 2$.
@@
I have demonstrated that thermal disorder of spins can frustrate the relative motion of dopants and lead to a dynamical pairing mechanism. The effect is closely related to the thermally induced localization phenomenon recently discovered for single holes Nielsen (2024a, b). By carefully analyzing its origin and behavior at infinite temperature, it is, furthermore, clear that the effect persists for all values of the spin coupling versus the hopping. For diminishing temperatures, ferromagnetic interaction is expected to lead to a crossover to delocalized and uncorrelated holes. For antiferromagnetic interactions, on the other hand, a crossover from fluctuation induced pairing to confinement of the holes emerges at low temperatures Nielsen (2023). Moreover, it demonstrates that the localization phenomenon Nielsen (2024a, b) is not sensitive to further doping. The results finally show that it is feasible to experimentally test my predictions using currently available techniques with ultracold atoms in optical lattices. One could also engineer similar models using Rydberg arrays Schauss et al. (2015); Labuhn et al. (2016); Bernien et al. (2017); Guardado-Sanchez et al. (2018); Lienhard et al. (2018) or polar molecules Gorshkov et al. (2011). In the future, the generality of the discovered effect may be tested by, e.g., tuning in trans-leg hopping or spin superexchange. The robustness of the phenomenon to such alterations is presently unclear. The nature of the effect is, however, very general – it comes down to a randomly fluctuating field emerging as the dopants separate, frustrating their relative motion.
@@ Acknowledgements.
The author thanks Pascal Weckesser, Johannes Zeiher, Pavel Kos, and J. Ignacio Cirac for helpful input and discussions. A special thanks goes to Pascal Weckesser for computing the soft core potentials. This work has been supported by the Carlsberg Foundation through a Carlsberg Internationalisation Fellowship, Grant No. CF21_0410.

### v3

@@ Abstract
Figure 3: Pair subdiffusion and steady state. (a) For $J_{\parallel}=0$, the hole pair spreads subdiffusively, $X_{\rm rms}\sim\tau^{\nu}$ at long times [top] with an exponent $\nu$ decreasing with $J_{\perp}/t$ [bottom]. Black- and gray-filled lines [top] give the dynamics for the full [Eq. (1)] and effective models, with corresponding black- and gray-enscribed points for the fitted exponents [bottom]. (b) Steady state rms distances $l_{X}$ and $l_{\Delta}$ of the center-of-mass, $X$, and relative, $\Delta$, coordinates of the holes as a function of $J_{\perp}/t$ for specified $J_{\parallel}$. For $J_{\parallel}=0$, $l_{X}$ saturates at $\sim\sqrt{N}$. (c) Correlation coefficient ${\rm corr}(x_{1},x_{2})=\braket{x_{1}x_{2}}/[\sigma(x_{1})\sigma(x_{2})]$, and (d) relative hole-hole correlator [Eq. (10)] for holes on the same rung ($\Delta=0$) vs $J_{\perp}/t$. (e) Relative hole-hole correlator vs $|\Delta|$ for indicated values of $J_{\perp}/t$ and $J_{\parallel}$.
Eventually, the system settles into a steady state, such that $X_{\rm rms}\to l_{X},\Delta_{\rm rms}\to l_{\Delta}$ [Fig. 3(b)]. The associated asymptote of the correlation coefficient [Fig. 3(c)] shows attractive correlations over the entire investigated range. Moreover, the relative hole-hole correlator in Figs. 3(d) and 3(e) show that the probability of finding holes in close proximity can be up to an order of magnitude larger than expected from their marginals. For $J_{\perp}=J_{\parallel}\geq 3t$, there are minor finite-size corrections, whilst smaller spin couplings results in more spread out holes and significant corrections [31]. Indeed, for $J_{\parallel}=0$, the subdiffusive behavior leads to a system-size dependent steady state, and I find $l_{X}\sim\sqrt{N}$. In turn, the correlation coefficient [Fig. 3(b)] behaves as $1-{\rm corr}(x_{1},x_{2})\propto l_{\Delta}^{2}/N$. The relative hole-hole correlator is, however, found to be well-converged for any $J_{\perp}>3t$ [31]. For $J_{\perp}\leq 3t$ and both values of $J_{\parallel}$, there are substantial finite-size effects that suppress the pairing signals, because it foremost restricts the center-of-mass motion. In this manner, pairing is demonstrating across the entire investigated regime and persists for $J_{\perp},J_{\parallel}<t$.
The origin of this dynamical pairing is exactly the same as for the thermal localization phenomenon discovered recently for the motion of single holes [28, 29]. Indeed, because the induced interaction $V_{\boldsymbol{\sigma}}(x_{1},x_{2})$ inevitably fluctuates to large values in the relative distance $2|\Delta|=|x_{1}-x_{2}|$ [Fig. 2(b)], the holes, at the very least, back-reflect once they meet their classical turning points, corresponding to emergent Anderson localization [32] for strong disorder [33]. For nonzero $J_{\parallel}$, the holes remain localized – a finite $l_{X}$, because even when they are on the same rung, they will encounter a fluctuating potential taking on values in the interval $[-|J_{\parallel}|,+|J_{\parallel}]$ [Fig. 2(c)], leading to Anderson localization for weak disorder [33].
@@
Quantum simulators based on Rydberg-dressed atoms in optical lattices natively realizes spin-specific density interactions [34]
${\hat{H}}_{J}=\sum_{{\bf i}\neq{\bf j}}\frac{J_{{\bf i}-{\bf j}}}{2}\hat{n}_{\uparrow{\bf i}}\hat{n}_{\uparrow{\bf j}}$ (11)
@@
of the internal atomic state $\ket{\uparrow}$, by optically dressing it with a higher-lying Rydberg state. Here, $J_{{\bf i}-{\bf j}}$ takes on a soft core shape [35, 31]. Such systems were recently demonstrated to enter the itinerant regime [36]. Finally, a Feshbach resonance used to enhance the onsite interaction deep into the Mott-insulating phase ensures at most one spin per site, negligible superexchange, and hereby realizes the desired itinerant spin model governed by ${\hat{H}}_{J}$ in Eq. (11) and the constrained hopping ${\hat{H}}_{t}$ in Eq. (1). Precise control of the temperature in such systems is difficult, but the hole dynamics may still be investigated in the following manner [28]. (1) initialize holes at $x_{1}=x_{2}=0$ by applying a strong repulsive light field to those sites [6], (2) rotate all spins from the fully polarized spin sample in the non-interacting $\ket{\downarrow}$ state into the equal superposition state |$\ket{\Psi_{0}}=\prod_{{\bf i}\neq(0,0),(0,1)}\frac{1}{\sqrt{2}}\left[\hat{c}^{\dagger}_{{\bf i}\uparrow}+\hat{c}^{\dagger}_{{\bf i}\downarrow}\right]\ket{0}$, and (3) turn off the focused light fields, freeing the holes to move along the ladder [12].
@@
This initially seems distinct from the infinite temperature scenario above. The hole-hole correlator dynamics, $C_{hh}(x_{1},x_{2};\tau)=2^{-N}\sum_{{\boldsymbol{\sigma}}}|a_{\boldsymbol{\sigma}}(x_{1},x_{2};\tau)|^{2}$, is however unchanged, whereby the holes move as at infinite temperature. The polarized Ising form of the interaction is crucial for this exact equivalence, and one should note that spin correlators would presumably differ from their infinite temperature counterpart. The computation now follows the same recipe as before. The resulting correlation coefficient dynamics [Eq. (9)] and relative correlation function [Eq. (10)] are given in Figs. 4(a) and 4(b), respectively. This is averaged over $N_{S}=1000$ realizations and given in two instances [inset of Fig. 4(a)]: (1) isotropic couplings akin to the $J_{\parallel}=J_{\perp}$ case, and (2) anisotropic couplings where the $30P_{3/2}$ orbital of the Rydberg state is rotated to lie along the rungs using an external magnetic field [34], suppressing coupling in other directions similar to low values of $J_{\parallel}$ [31]. Crucially, we observe that the appearing disorder-induced attraction between the holes can be probed for experimentally feasible system sizes and timescales. Indeed, stroboscopic dressing [36] makes $20$ tunneling events feasible, while longer timescales are currently inhibited by the intrinsic lifetime of the Rydberg state [37].
@@
Figure 4: Rydberg dressing. (a) Dynamics of the correlation coefficient for isotropic (orange) and strong rung (purple) couplings. (b) Relative correlator for holes on the same rung, $\Delta=0$. Dashed lines: steady state values. Nearest neighbor rung interaction: $J_{{\bf e}_{y}}=8t$. Insets: $30P_{3/2}$ orbital orientation dictated by the magnetic fields, ${\bf B}$. System size: $N=21\times 2$.
@@
I have demonstrated that thermal spin disorder can frustrate the relative motion of dopants and lead to dynamical pairing. The nature of the effect is very general: it comes down to a randomly fluctuating field emerging as the dopants separate. It is closely related to the thermally induced localization phenomenon recently discovered for single holes [28, 29]. By carefully analyzing its origin and behavior at infinite temperature, it is clear that the effect persists for all values of the spin coupling versus the hopping. For diminishing temperatures, ferromagnetic interactions should lead to a crossover to delocalized and uncorrelated holes. For antiferromagnetic interactions, a crossover from fluctuation induced pairing to confinement of the holes emerges at low temperatures [19]. As demonstrated, my predictions are feasibly testable using available techniques with ultracold atoms in optical lattices. Alternative platforms include Rydberg arrays [38, 39, 40, 41, 42] and polar molecules [43]. The generality of the discovered effect may be tested by adding trans-leg hopping, $t_{\perp}$, or spin superexchange $\propto\alpha J$. I expect pairing to be robust as long as the hopping along the ladder dynamically dominates the perturbations, $t\gg t_{\perp},\alpha J$. Such computations go beyond the presently described methodology, but may be amenable to DMRG approaches [44]. This will be addressed in future work.
@@ Acknowledgements.
The author thanks Pascal Weckesser, Johannes Zeiher, Pavel Kos, and J. Ignacio Cirac for helpful input and discussions. Special thanks go to Pascal Weckesser for computing the soft core potentials. This work has been supported by the Carlsberg Foundation through a Carlsberg Internationalisation Fellowship, Grant No. CF21_0410.
The data that support the findings of this article are openly available [45].